# RC-positivity, comparison theorems and prescribed Hermitian-Yang-Mills tensors I

### 13 Mar 2026 v2

@@ Abstract.
As applications, we obtain quantitative Chern number inequalities applicable to both holomorphic vector bundles and Fano manifolds. The proof is based on a new comparison theorem for Hermitian-Yang-Mills tensors.
@@ 1. Introduction
By Serre duality, the corresponding conjugate equation of (1.4) admits a solution when $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is (quasi-)negative definite; see Remark 5.2 for further details. The method for solving equation (1.4) differs significantly from those for equations (1.1) and (1.2), and we provide a brief overview of the key ideas involved. The uniqueness part of Theorem 1.1 is established by the following comparison theorem for Hermitian-Yang-Mills tensors of Hermitian vector bundles:
@@
By combining Theorem 1.1 and Theorem 1.2, we derive a quantitative version of Chern number inequalities, extending those associated with Hermitian-Einstein metrics.
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
(1.11) $a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0},$
@@
(1.12) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or $M$ is projective.
@@ 5. Proofs of main results
We demonstrate that Theorem 1.1 holds in a slightly more general setting.
@@
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is quasi-positive definite. Then for any Hermitian positive definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
@@
Recall that a tensor is quasi-positive if it is nonngeative everywhere and positive definite at some point. Actually Theorem 5.8 follows from Theorem 1.1 and the following result.
Proposition 5.9.
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a smooth Hermitian metric $h_{0}$ on $E$ such that the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is quasi-positive definite. Then there exists a smooth Hermitian metric $h_{1}$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{1}}$ is positive definite.
@@
We employ the strategy utilized in the proof of Theorem 5.6. We fix the center point $h_{0}\in\mathrm{Herm}^{+}(E)$ and so $F(h_{0})=\mathrm{Id}_{E}$. We show that $G$ is invertible around $h_{0}$. In particular, when $\varepsilon>0$ is small, there exists some $h_{1}$ on $E$ such that
(5.84) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{1}}=\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}+\varepsilon h_{0}>0.$
Indeed, for any $\Psi\in\mathrm{Herm}(E,h_{0})$,
(5.85) $\mathscr{L}_{\mathrm{Id}_{E}}(\Psi)=\Delta_{{\partial}^{h_{0}}}\Psi+\Omega_{0}\cdot\Psi.$
where $\Omega_{0}=S^{h_{0}}\in\Gamma(M,E^{*}\otimes E)$. To apply the framework established in Theorem 5.6, it suffices to demonstrate the existence of a constant $C=C(M,\omega_{g},h_{0})$ such that for each $\Psi\in W^{1,2}(M,E^{*}\otimes E)$,
(5.86) $(\mathscr{L}_{\mathrm{Id}_{E}}\Psi,\Psi)\geq C(\Psi,\Psi),$
and this implies
(5.87) $(\mathscr{L}_{\mathrm{Id}_{E}}\Psi,\Psi)\geq\min\left\{\frac{C}{2},\frac{1}{2}\right\}\|\Psi\|_{W^{1,2}(M,E^{*}\otimes E)}.$
If it is not valid, for any $m\in\mathbb{N}$, there exists nonzero $\Psi_{m}\in W^{1,2}(M,E^{*}\otimes E)$ such that
(5.88) $({\partial}^{h_{0}}\Psi_{m},{\partial}^{h_{0}}\Psi_{m})+(\Omega\cdot\Psi_{m},\Psi_{m})<\frac{1}{m}(\Psi_{m},\Psi_{m}).$
We use the normalization trick and define
(5.89) $\Phi_{m}:=\frac{\Psi_{m}}{\left\|\Psi_{m}\right\|_{L^{2}(M,E^{*}\otimes E)}}.$
One has $\left\|\Phi_{m}\right\|_{L^{2}(M,E^{*}\otimes E)}=1$ and
(5.90) $({\partial}^{h_{0}}\Phi_{m},{\partial}^{h_{0}}\Phi_{m})+(\Omega\cdot\Phi_{m},\Phi_{m})<\frac{1}{m}(\Phi_{m},\Phi_{m})=\frac{1}{m}.$
By (5.90) and the normalization condition, there exists $C=C(M,\omega_{g})$ such that,
(5.91) $\left\|\Phi_{m}\right\|_{W^{1,2}(M,E^{*}\otimes E)}\leq C.$
By the weak compactness theorem and the compact embedding of $W^{1,2}(M,E^{*}\otimes E)\subset L^{2}(M,E^{*}\otimes E)$, there exists a subsequence $\{\Phi_{m_{k}}\}$ and $\Phi\in W^{1,2}(M,E^{*}\otimes E)$ such that $\Phi_{m_{k}}\rightarrow\Phi$ in the weak $W^{1,2}(M,E^{*}\otimes E)$ sense and $\Phi_{m_{k}}\rightarrow\Phi$ in $L^{2}(M,E^{*}\otimes E)$ sense. It is clear that $\left\|\Phi\right\|_{L^{2}(M,E^{*}\otimes E)}=1$. By the weak convergence theorem,
[…]
Therefore, we obtain that ${\partial}^{h_{0}}\Phi=0$. On the other hand, by the $L^{2}$-convergence,
(5.93) $(\Omega\cdot\Phi,\Phi)=\lim_{m_{k}\rightarrow\infty}(\Omega\cdot\Phi_{m_{k}},\Phi_{m_{k}})\leq\lim_{m_{k}\rightarrow\infty}\frac{1}{m_{k}}=0.$
Since $\Omega$ is quasi-positive, there exists an open set $U\subset M$ such that $\Phi\equiv 0$ on $U$. Since ${\partial}^{h_{0}}\Phi=0$, $\Phi\equiv 0$ on $M$ which contradicts to the fact $\left\|\Phi\right\|_{L^{2}(M,E^{*}\otimes E)}=1$. ∎
@@
Proposition 5.9 provides an alternative proof of the main results in [Yang24] and [LL22].
@@ 6. Chern number inequalities
In this section, we prove Theorem 1.6 in a slightly general setting:
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ satisfying
(6.1) $a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0},$
@@
(6.2) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n},$
if $ab>0$ or there exists a holomorphic line bundle $L$ on $M$ satisfying $\int_{M}c_{1}(L)\wedge\omega_{g}^{n}\neq 0$.
@@
We assume that $b\geq a>0$. By Theorem 1.1, there exists a smooth Hermitian metric $h$ on $E$ such that
(6.3) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}.$
@@
Since $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}$, we get
(6.8) $s_{h}=\mathrm{tr}_{h}h_{0},\quad\mbox{ and}\quad\left|\mathrm{Ric}^{(2)}\right|^{2}=|h_{0}|_{h}^{2}.$
Therefore, we obtain
$\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(r|T|^{2}+\frac{n-1}{n}\left((\mathrm{tr}_{h}h_{0})^{2}-r|h_{0}|^{2}_{h}\right)\right)\omega_{g}^{n}.$
Since $0<a\cdot h_{0}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot h_{0}$ and $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=h_{0}$, we deduce that
(6.9) $0<a\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\leq b\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h}.$
By Theorem 1.2, we obtain
(6.10) $a\cdot h\leq h_{0}\leq b\cdot h.$
In particular, the following estimates hold
(6.11) $(\mathrm{tr}_{h}h_{0})^{2}-r|h_{0}|^{2}_{h}\geq-\frac{r(r-1)}{2}(b-a)^{2}\geq-r^{2}(b^{2}-a^{2}).$
@@
For the case $a\leq b<0$, we know that
(6.12) $0<-b\cdot h_{0}^{*}\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}^{*}}\leq-a\cdot h_{0}^{*}$
as Hermitian tensors in $\Gamma(M,E\otimes\overline{E})$ under the isomorphism $E\cong(E^{*})^{*}$. Hence, there exists a Hermitian metric $h^{*}$ on $E^{*}$ such that
(6.13) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}}=h^{*}_{0}.$
Moreover, one has
(6.14) $0<-b\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}}\leq-\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{*}_{0}}\leq-a\cdot\sqrt{-1}\Lambda_{\omega_{g}}R^{h^{*}},$
and so $bh^{*}\leq h_{0}^{*}\leq-ah^{*}$. This implies
(6.15) $\int_{M}\left(2rc_{2}(E^{*})-(r-1)c^{2}_{1}(E^{*})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
Since $c_{1}^{2}(E^{*})=c_{1}^{2}(E)$ and $c_{2}(E^{*})=c_{2}(E)$, we get the result.
Suppose that there exists a holomorphic line bundle $L$ over $M$ with $\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}>0$. It is easy to show that there exists a smooth Hermitian metric $h^{L}$ on $L$ such that
(6.16) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h^{L}}=c_{0}=\frac{n\int_{M}c_{1}(L)\wedge\omega_{g}^{n-1}}{\int_{M}\omega_{g}^{n}}>0.$
We consider the twisted Hermitian vector bundle $E_{k}:=E\otimes L^{k}$. The Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ is given by
(6.17) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}\right)\otimes h^{kL}+h_{0}\otimes\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h^{kL}}\right).$
In particular,
(6.18) $\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)=\left(\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}+kc_{0}h_{0}\right)\otimes h^{kL}.$
We fix a positive integer $k$ such that $a+kc_{0}>0$. Then the Hermitian-Yang-Mills tensor of $(E_{k},h_{0}\otimes h^{kL})$ satisfies
(6.19) $0<(a+kc_{0})(h_{0}\otimes h^{kL})\leq\Lambda_{\omega_{g}}\left(\sqrt{-1}R^{h_{0}\otimes h^{kL}}\right)\leq(b+kc_{0})(h_{0}\otimes h^{kL}).$
In particular, by the first case, we deduce that
(6.20) $\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}\geq-\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
On the other hand, it is well-known that
$\int_{M}\left(2rc_{2}(E\otimes L^{k})-(r-1)c^{2}_{1}(E\otimes L^{k})\right)\wedge\omega_{g}^{n-2}=\int_{M}\left(2rc_{2}(E)-(r-1)c^{2}_{1}(E)\right)\wedge\omega_{g}^{n-2}.$
@@
It is expected that the projectivity condition in Theorem 1.6 is redundant, i.e., the Chern number inequality (1.12) holds for any $a,b\in{\mathbb{R}}$ satisfying (1.11).
Corollary 6.3.
Let $M$ be a Fano manifold. If there exists a Kähler metric $\omega_{g}$ such that
(6.21) $a\omega_{g}\leq\mathrm{Ric}(\omega_{g})\leq b\omega_{g}$
for some constants $a>0,b>0$, then the following Chern number inequality holds
(6.22) $\int_{M}\left((n-1)c^{2}_{1}(M)-2nc_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(n-1)\left(b-a\right)^{2}}{8\pi^{2}n}\int_{M}\omega_{g}^{n}.$
Remark 6.4.
In the proof of Corollary 6.3, we use the existence of a Hermitian metric $h$ on $M$ satisfying
(6.23) $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}=g,$
where $g$ is a Kähler metric. This condition enriches the geometric interplay between $h$ and $g$, allowing for sharper control in the curvature estimates. In particular, when $h$ is a Kähler or balanced metric– i.e., $d\omega^{n-1}_{h}=0$–the associated Chern number inequality can be strengthened to
(6.24) $\int_{M}\left(nc^{2}_{1}(M)-2(n+1)c_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(n^{2}-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
This refines and extends Yau’s classical Chern number inequalities [Yau77] for Fano Kähler-Einstein manifolds.

### 30 Mar 2026 v3

@@ Abstract.
The proof is based on a new comparison theorem for Hermitian-Yang-Mills tensors. Inspired by these results, we have also derived quantitative Chern number inequalities that apply to both holomorphic vector bundles and compact Kähler manifolds.
@@ 1. Introduction
The positivity of $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ can be weakened to the positivity of the integral of the minimum eigenvalue of $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ over the manifold (Theorem 5.8). Moreover, by Serre duality, the corresponding conjugate equation of (1.4) admits a solution when $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}$ is negative definite; see Remark 5.2 for further details. The method for solving equation (1.4) differs significantly from those for equations (1.1) and (1.2), and we provide a brief overview of the key ideas involved. The uniqueness part of Theorem 1.1 is established by the following comparison theorem for Hermitian-Yang-Mills tensors of Hermitian vector bundles:
@@
As inspired by Theorem 1.1 and the comparsion result described in Theorem 1.2, we derive a quantitative version of Chern number inequalities, extending those associated with Hermitian-Einstein metrics.
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h$ on $E$ satisfying
(1.11) $a\cdot h\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq b\cdot h,$
@@
(1.12) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
Moreover, we establish the following Chern number inequalities extending those associated with Kähler-Einstein metrics ([Yau77]).
Theorem 1.7.
Let $(M,\omega_{g})$ be a compact Kähler manifold with $\dim M\geq 2$. If there exist two constants $a,b\in\mathbb{R}$ such that
(1.13) $a\omega_{g}\leq\mathrm{Ric}(\omega_{g})\leq b\omega_{g},$
then the following Chern number inquality holds
(1.14) $\int_{M}\left(nc^{2}_{1}(M)-2(n+1)c_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(n^{2}-2)(b-a)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
@@ 5. Proofs of main results
We demonstrate that Theorem 1.1 holds in a more general setting. Let $h$ be a smooth Hermitian metric on $E$. The minimal eigenvalue function $\kappa_{h}:M\rightarrow\mathbb{R}$ of the Hermitian-Yang-Mills tensor $S^{h}$ is defined as
(5.83) $\kappa_{h}(x)=\inf_{0\neq v\in E_{x}}\frac{h(S^{h}v,v)}{h(v,v)}.$
It is easy to see that $S^{h}>0$ with respect to $h$ if and only if $\kappa_{h}(x)>0$ for any $x\in M$.
@@
Let $E$ be a holomorphic vector bundle over a compact Kähler manifold $(M,\omega_{g})$. Suppose that there exists a Hermitian metric $h_{0}$ on $E$ such that
(5.84) $\int_{M}\kappa_{h_{0}}(x)\omega_{g}^{n}>0.$
Then for any Hermitian positive definite tensor $P\in\Gamma(M,E^{*}\otimes\overline{E}^{*})$, there exists a unique smooth Hermitian metric $h$ on $E$ such that
@@
By using the condition (5.84), there exists some $\lambda_{0}>0$ such that
(5.86) $\int_{M}\kappa_{h_{0}}(x)\omega_{g}^{n}=\lambda_{0}\int_{M}\omega_{g}^{n}.$
Since $\kappa_{h_{0}}$ is Lipschitz, by Hodge theory, there exists a function $f\in C^{2,\alpha}(M,\mathbb{R})$ such that
(5.87) $\kappa_{h_{0}}+\Delta_{g}f=\lambda_{0}.$
We define a new metric $\widetilde{h}=e^{-f}h$. By (3.6), one has
(5.88) $\Lambda_{\omega_{g}}\sqrt{-1}R^{\widetilde{h}}-\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{0}}={\partial}^{*,h_{0}}\left(({\partial}^{h_{0}}H)\cdot H^{-1}\right)$
where $H=\widetilde{h}\cdot h_{0}^{-1}=e^{-f}\cdot\mathrm{Id}_{E}$. A straightforward computation shows
${\partial}^{*,h_{0}}\left(({\partial}^{h_{0}}H)\cdot H^{-1}\right)$ $=$ ${\partial}^{*,h_{0}}\left(-\left(e^{-f}{\partial}f\otimes\mathrm{Id}_{E}\right)\cdot H^{-1}\right)$
$=$ $\sqrt{-1}\Lambda_{\omega_{g}}\overline{{\partial}}\left(-{\partial}f\otimes\mathrm{Id}_{E}\right)$
$=$ $-\sqrt{-1}\Lambda_{\omega_{g}}\overline{{\partial}}{\partial}f\cdot\mathrm{Id}_{E}=\Delta_{g}f\cdot\mathrm{Id}_{E}.$
Moreover, one has
$\kappa_{\widetilde{h}}(x)=\inf_{0\neq v\in E_{x}}\frac{{\widetilde{h}}(S^{\widetilde{h}}v,v)}{{\widetilde{h}}(v,v)}$ $=$ $\inf_{0\neq v\in E_{x}}\frac{{\widetilde{h}}(S^{h_{0}}v,v)}{{\widetilde{h}}(v,v)}+\Delta_{g}f(x)$
$=$ $\inf_{0\neq v\in E_{x}}\frac{h_{0}(S^{h_{0}}v,v)}{h_{0}(v,v)}+\Delta_{g}f(x)$
$=$ $\kappa_{h_{0}}(x)+\Delta_{g}f(x)=\lambda_{0}.$
In particular, the Hermitian-Yang-Mills tensor $\Lambda_{\omega_{g}}\sqrt{-1}R^{\widetilde{h}}$ is positive definite with respect to $\widetilde{h}$. By a perturbation trick, one can find a smooth Hermitian metric $h_{1}$ in a small neighborhood of $\widetilde{h}$ such that $\Lambda_{\omega_{g}}\sqrt{-1}R^{h_{1}}$ is positive definite with respect to $h_{1}$. Now Theorem 5.8 follows by Theorem 1.1. ∎
@@
Theorem 5.8 also holds on general Hermitian manifolds and provides an alternative proof of the main results in [Yang24] and [LL22].
@@ 6. Chern number inequalities
In this section, we prove Theorem 1.6 and Theorem 1.7.
@@
Let $(M,\omega_{g})$ be a compact Kähler manifold and $E$ be a holomorphic vector bundle of rank $r$ over $M$. Suppose that there exists a Hermitian metric $h$ on $E$ satisfying
(6.1) $a\cdot h\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq b\cdot h,$
@@
(6.2) $\int_{M}\left((r-1)c^{2}_{1}(E)-2rc_{2}(E)\right)\wedge\omega_{g}^{n-2}\leq\frac{r(r-1)\left(b-a\right)^{2}}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
@@
Let $\{\lambda_{1},\cdots,\lambda_{r}\}$ be the eigenvalues of $\Lambda_{\omega_{g}}\sqrt{-1}R^{h}$ with respect to $h$. Since $ah\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq bh$, it is obvious that
(6.7) $a\leq\lambda_{k}\leq b,$
and so
(6.8) $s_{h}^{2}-r\left|\mathrm{Ric}^{(2)}\right|^{2}=\left(\sum_{k=1}^{r}\lambda_{k}\right)^{2}-r\sum_{k=1}^{r}\lambda_{k}^{2}=-\sum_{i<j}(\lambda_{i}-\lambda_{j})^{2}\geq-\frac{r(r-1)}{2}(a-b)^{2}.$
@@
The following result refines and extends Yau’s classical Chern number inequalities [Yau77] for Kähler-Einstein manifolds.
Theorem 6.2.
Let $(M,\omega_{g})$ be a compact Kähler manifold with $\dim M\geq 2$. If there exist two constants $a,b\in\mathbb{R}$ such that
(6.9) $a\omega_{g}\leq\mathrm{Ric}(\omega_{g})\leq b\omega_{g},$
then the following Chern number inquality holds
(6.10) $\int_{M}\left(nc^{2}_{1}(M)-2(n+1)c_{2}(M)\right)\wedge\omega_{g}^{n-2}\leq\frac{(a-b)^{2}(n^{2}-2)}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
Proof.
It is known that
(6.11) $\int_{M}c^{2}_{1}(M)\wedge\omega_{g}^{n-2}=\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(s^{2}-\left|\mathrm{Ric}\right|^{2}\right)\omega_{g}^{n},$
and
(6.12) $\int_{M}c_{2}(M)\wedge\omega_{g}^{n-2}=\frac{1}{8\pi^{2}n(n-1)}\int_{M}\left(s^{2}-2\left|\mathrm{Ric}\right|^{2}+|R|^{2}\right)\omega_{g}^{n},$
where $R$ is the curvature tensor, $\mathrm{Ric}$ is the Ricci curvature, and $s$ is the scalar curvature of the Kähler metric $g$. Let $T\in\Omega^{1,1}(M,T^{*1,0}M\otimes T^{*0,1}M)$ be the tensor given by
(6.13) $T_{i\overline{j}k\overline{\ell}}=R_{i\overline{j}k\overline{\ell}}-\frac{1}{n}g_{i\overline{j}}R_{k\overline{\ell}}-\frac{1}{n}R_{i\overline{j}}g_{k\overline{\ell}}-\frac{1}{n(n+1)}g_{i\overline{\ell}}g_{k\overline{j}}s+\frac{n+2}{n^{2}(n+1)}g_{i\overline{j}}g_{k\overline{\ell}}s.$
A straightforward computation shows that
(6.14) $|T|^{2}=|R|^{2}-\frac{2}{n}\left|\mathrm{Ric}\right|^{2}+\frac{2}{n^{2}(n+1)}s^{2}.$
By (6.11), (6.12) and (6.14), one has
$\int_{M}\left(2(n+1)c_{2}(M)-nc_{1}(M)^{2}\right)\wedge\omega_{g}^{n-2}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left(s^{2}-(n+2)\left|\mathrm{Ric}\right|^{2}+(n+1)|R|^{2}\right)\omega_{g}^{n}$
$=$ $\frac{1}{4\pi^{2}n(n-1)}\int_{M}\left((n+1)|T|^{2}+\left(1-\frac{2}{n^{2}}\right)\left(s^{2}-n\left|\mathrm{Ric}\right|^{2}\right)\right)\omega_{g}^{n}.$
Let $\{\lambda_{1},\cdots,\lambda_{n}\}$ be the eigenvalues of $\mathrm{Ric}(\omega_{g})$ with respect to $g$. Since $a\omega_{g}\leq\mathrm{Ric}(\omega_{g})\leq b\omega_{g}$, it is obvious that $a\leq\lambda_{k}\leq b$, and so
(6.15) $s^{2}-n\left|\mathrm{Ric}\right|^{2}=\left(\sum_{k=1}^{n}\lambda_{k}\right)^{2}-n\sum_{k=1}^{n}\lambda_{k}^{2}=-\sum_{i<j}(\lambda_{i}-\lambda_{j})^{2}\geq-\frac{n(n-1)}{2}(a-b)^{2}.$
Therefore, we obtain
$\int_{M}\left(2(n+1)c_{2}(M)-nc_{1}(M)^{2}\right)\wedge\omega_{g}^{n-2}$ $\geq$ $\frac{n^{2}-2}{4\pi^{2}n^{3}(n-1)}\int_{M}\left(s^{2}-n\left|\mathrm{Ric}\right|^{2}\right)\omega_{g}^{n}$
$\geq$ $-\frac{(a-b)^{2}(n^{2}-2)}{8\pi^{2}n^{2}}\int_{M}\omega_{g}^{n}.$
This completes the proof. ∎
@@
When $h$ is a Kähler or balanced metric on a compact complex manifold– i.e., $d\omega^{n-1}_{h}=0$– and satisfies the inequality
(6.16) $a\cdot h\leq\Lambda_{\omega_{g}}\sqrt{-1}R^{h}\leq b\cdot h,$
for some constants $a,b\in{\mathbb{R}}$, then analogous Chern number inequalities to those in (6.10) hold.